For each of the data sets, create a scatter plot with the regression line and answer these questions. What is r, r2, Critical Value (CV)? In general, what does r2 mean? What does r2 mean specifically for the data set? What is the regression equation? Is there any positive or negative correlation? Interpret b1. Interpret b0. Forecast the next year.

Year

Cost (Thousands of dollars)

1

100

2

99

3

102

4

80

5

90

6

82

7

80

8

76

9

85

10

70
Answer #1

Scatterplot 120 100 0 4 6 8 10 14 X(year)

X(year) Y(cost) (x-x̅)² (y-ȳ)² (x-x̅)(y-ȳ)
1 100 20.25 184.96 -61.2
2 99 12.25 158.76 -44.1
3 102 6.25 243.36 -39
4 80 2.25 40.96 9.6
5 90 0.25 12.96 -1.8
6 82 0.25 19.36 -2.2
7 80 2.25 40.96 -9.6
8 76 6.25 108.16 -26
9 85 12.25 1.96 -4.9
10 70 20.25 268.96 -73.8
ΣX ΣY Σ(x-x̅)² Σ(y-ȳ)² Σ(x-x̅)(y-ȳ)
total sum 55 864 82.5 1080.4 -253
mean 5.5 86.4 SSxx SSyy SSxy

sample size ,   n =   10          
here, x̅ =   5.5       ȳ =   86.4  
                  
SSxx =    Σ(x-x̅)² =    82.5          
SSxy=   Σ(x-x̅)(y-ȳ) =   -253          
                  
slope ,    ß1 = SSxy/SSxx =   -3.066666667          
                  
intercept,   ß0 = y̅-ß1* x̄ =   103.2666667          
                  
so, regression line is   Ŷ =   103.2667   +   -3.0667   *x
                  
SSE=   (Sx*Sy - S²xy)/Sx =    304.53          
                  
std error ,Se =    √(SSE/(n-2)) =    6.1698          
                  
correlation coefficient ,    r = Sxy/√(Sx.Sy) =   -0.8474          
                  
R² =    (Sxy)²/(Sx.Sy) =    0.7181      

---------------------------

r = correlation = -0.8474

R² = coefficient of determination = 0.7181

in general, R² represents how much percentage of variation in observation of y is explained by variable X

--------------------------------

What does r2 mean specifically for the data set here,

Here,71.81 % variation in observations of cost is explained by variable year

-----------------------------

What is the regression equation?

from above calculations

regression eqn is

Ŷ =   103.2667    -3.0667   *x

-------------------------------------

Is there any positive or negative correlation

since, the value of r is negative,so correlation is negative

--------------------------

Interpret b1. Interpret b0.

b1--> for 1 unit increase in Year,the cost will get decrease by 3.0667(thousand of dollars)

bo--> when year=0,then cost will be 103.2667(thousand of dollars)

-----------------------

Forecast the next year

Ŷ =   103.2667    -3.0667   *x

for x=11

Ŷ =   103.2667    -3.0667   *11 = 69.533 (thousand of dollars)